a 65 mm diameter solid shaft is to be welded to a flat plate by a fillet weld around the circumference of the shaft. determine the size of the weld if the torque on the shaft is 3 kn m. the allowable shear stress in the weld is 70 mpa.


1
Expert's answer
2022-02-23T00:40:01-0500

"\tau_{max}=\frac{2.83T}{πsd^2}"

"s=\frac{2.83T}{π \tau_{max}d^2}"

Where s is the weld size

"s=\frac{2.83*3*10^3}{π*70*10^6*65^2*10^{-2}}"

="9.13*10^{-3}m"

=9.13mm

Hence weld size =10mm